# Max-Rank: Efficient Multiple Testing for Conformal Prediction

### v3

@@ 3.1 The max-rank algorithm
Consider running $m$ hypothesis tests in parallel, and for any particular testing dimension $k\in[m]$ the statistics forming an empirical null distribution are given by $\textbf{s}_{k}\in\mathbb{R}^{n}$. In the conformal setting, $\mathbf{s}_{k}$ equates the set of nonconformity scores on $\mathcal{D}_{cal}$. We now desire finding adjusted significance levels $\alpha_{k}$ for all dimensions such that the FWER is globally controlled at target level $\alpha$.
We algorithmically describe our proposed max-rank correction in Algo. 2. First, the empirical nulls of equal size are collected in a matrix of test statistics $\mathbf{S}=[s_{i,k}]\in\mathbb{R}^{n\times m}$, where $i\in[n]$ and $k\in[m]$. The $m$ columns span the tested hypotheses, and the $n$ rows represent values in each null distribution. Next, we obtain a similar matrix containing the column-wise ranks of $\mathbf{S}$ as $\mathbf{R}=[r_{i,k}]\in\mathbb{N}^{n\times m}$, where also $i\in[n]$ and $k\in[m]$. By Eq. 4, the search for a significance level $\alpha_{k}$ for each test is equivalent to computing the conformal quantile $\hat{Q}(1-\alpha;\mathbf{s}_{k})$ in the space of test statistics. This in turn equates determining the value $s_{t,k}$, where $t=\hat{Q}(1-\alpha;\mathbf{r}_{k})$ denotes a quantile in rank space ensuring the desired error control (Remark 1). max-rank adjusts these rank quantiles for FWER control by forming a composite empirical null $\mathbf{r}_{\max}$ via the $l^{\infty}$-norm across testing dimensions, and using the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ to inform globally corrected significance levels for each hypothesis test $k\in[m]$.
@@
2:Output: Adjusted quantiles with FWER control at level $\alpha$ as $(\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m})\in\mathbb{R}^{m}$.
@@
6:Compute the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ as the $\lceil(n+1)(1-\alpha)/n\rceil$-th empirical quantile of $\mathbf{r}_{\max}$.
@@
8:Select the test statistic $s_{r_{\max},k}$ at rank $r_{\max}$ column-wise in $\mathbf{S}$ as the adjusted quantile $\hat{Q}_{1-\alpha,k}$ for every test dimension $k\in[m]$.
@@
Intuition. Our max-rank correction intuitively collapses multiple empirical nulls into a single ‘worst-case’ null for which the error-controlling quantile is computed. Thus, any potential dependencies across test statistics which are reflected in the rank ordering are efficiently aggregated in the ‘max-rank’ statistic $\mathbf{r}_{\max}$, and subsequently the selected rank quantile $r_{\max}$. By virtue of applying the $l^{\infty}$-norm, $r_{\max}$ ensures that adjusted significance levels provide FWER control.
If the empirical nulls are uncorrelated and the tested hypotheses can thus be considered independent, no benefit is obtained from max-rank’s aggregation step. We prove that in such cases our approach performs at least on par with the Bonferroni correction (§ 4), and never worse. On the other hand, if the tests are perfectly correlated each empirical null exhibits identical rank ordering, and max-rank will provide the optimal quantile solution for each individual testing dimension. For any cases with positively dependent tests inbetween the two extremes, our approach ensures superior performance over Bonferroni, and scales well even as the number of hypotheses grows (see § 5).
@@ 3.2 Relation to Westfall & Young corrections
A strong connection is apparent between max-rank and the max-T and min-P resampling-based corrections of Westfall and Young (1993), which offer similar FWER control. The max-T method provides adjusted p-values by comparing test statistics to the distribution of maximum statistics across hypotheses (Algo. 3), while the min-P method adjusts p-values based on the minimum observed p-values after resampling (Algo. 4). In both cases, existing dependency structure between tests is leveraged through repeated data resampling.
@@
Exploring these two corrections through the lens of our conformal prediction setting, we find that max-rank can be cast as a procedure that (1) naturally extends max-T to the rank space, and (2) describes a particular instance of min-P arising from our problem setting. As such, we argue that max-rank exhibits the advantages of both methods by (1) permitting its direct application to obtained test statistics as in max-T, whilst (2) remaining scale-invariant as in min-P. We next relate to each correction in more detail.
@@
Relation to max-T. Our conformal setting directly provides empirical null distributions for every hypothesis by use of the calibration set. The computed nonconformity scores on $\mathcal{D}_{cal}$ leverage available labels to produce statistics satisfying the tested null $H_{0}:Y_{n+1}=y$, and are thus interpretable as an analog to the required “data resampling under the null” (Westfall and Young, 1993). It follows that the adjusted p-values in Algo. 3 match the conformal p-values in Eq. 3 simply by replacing the score set $S$ with a maximum score set $S_{\max}$, i.e., replacing $\hat{P}_{n+1}(y;S)$ with $\hat{P}_{n+1}(y;S_{\max})$ [^2] And correcting for $(n+1)$ samples instead of $n$.. Instead of adjusting the p-values as in Algo. 3 we can directly adjust the significance level $\alpha$, which by Eq. 4 translates to computing the quantile $\hat{Q}(1-\alpha;S_{\max})$ instead of $\hat{Q}(1-\alpha;S)$. Comparing this to max-rank, we notice that the difference lies in its direct application to the maximum statistics, rather than translating them into rank space first (i.e., using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$). This means that max-T is sensitive to scale and requires test statistics across hypotheses to be directly comparable. Interestingly, such a “maximum of scores” approach has been intuitively employed by researchers in conformal prediction, e.g., Andéol et al. (2023); Schuster et al. (2021); Fisch et al. (2022), but never been recognized as an application case of Westfall and Young (1993).
@@
Relation to min-P. By the same resampling interpretation as above, unadjusted p-values for each hypothesis in Algo. 4 can be computed via conformal p-values in Eq. 3. Using the rank definition in Eq. 5, a given p-value $\hat{P}_{n+1}(y;S)$ can be expressed in terms of score ranking as $((n+1)-\text{rank}(s_{n+1};S))/(n+1)$ [^3] Observe that the permutation-based p-value in Eq. 3 counts the fraction of values above the quantity of interest, whereas the rank counts values below it.. Thus, computing the minimum p-values for min-P can be interpreted as taking the maximum over inverted and normalized ranks. Using Eq. 4, we again translate adjusting the p-values as adjusting the significance level via $\hat{Q}(\alpha;\mathbf{p}_{\min})$ instead. Note that this quantile is now taken over (min.) p-values, and thus only permits testing in p-value space. In contrast, max-rank maintains the simplicity of operating directly on the statistics, facilitating the interpretation of testing decisions.
@@ 4 THEORETICAL RESULTS
We next briefly describe our theoretical results for max-rank, deferring details and proofs to Appendix A. The multiple testing problem is first rephrased in terms of multivariate copulas, and our two key results state that (1) max-rank provides a valid solution with FWER control, and (2) is assured to perform better than Bonferroni under positive dependency of tests.
Recasting the multiple testing problem (§ A.1). Consider an $m$-dimensional random variable $\mathbf{X}\in\mathbb{R}^{m}$ and denote its c.d.f. as $F_{\mathbf{X}}(\mathbf{x})$. Using the interpretation in Remark 1, determining adjusted significance levels $(\alpha_{1},\dots,\alpha_{m})$ with joint Type-I error control at level $\alpha$ is then the equivalent to finding a vector $\mathbf{x}_{*}\in\mathbb{R}^{m}$ satisfying $F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha$. By using Sklar’s Theorem (Sklar, 1973) to express the c.d.f. in terms of a copula $C_{\mathbf{X}}$, and defining a vector of probabilities $(q_{*,1},\dots,q_{*,m})=(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))\in[0,1]^{m}$ restricted to a fixed value $q_{*}=q_{*,1}=\dots=q_{*,m}$, we can restate the problem as the search for a solution
@@
Obtaining a solution via max-rank (§ A.2). For a given solution to Eq. 6 the desired quantiles with FWER control are then computed as $\mathbf{x}_{*}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. These are equal to the quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo. 2. Leveraging operations in the rank space, we show that max-rank provides such a solution $q_{\max}$. We state this in the following proposition:
@@ 5 SIMULATION STUDY
Global FWER control. We compare the adjusted significance levels produced by both max-rank and Bonferroni across varying correlation strengths and for $m\in\{1,3,5,10,20\}$ in Fig. 1. We observe two key drawbacks for Bonferroni. Firstly, the correction becomes increasingly conservative as correlation between tests increases, with the adjusted level $\hat{\alpha}$ tending towards zero. Secondly, the rate accelerates as the number of tests increases. In contrast, max-rank suffers from neither of these pathologic behaviours and is able to tightly maintain FWER control at target level $\alpha$, ensuring good performance independently of both correlation strength and test dimensionality (Prop. 1).
@@
[…]
@@ 6 APPLICATION TO CONFORMAL PREDICTION
We now elaborate on how our proposed max-rank correction is useful for the conformal prediction setting introduced in § 2.1. First, we illustrate how multiple testing issues arise whenever multiple CP procedures are run in parallel. We then theoretically motivate why max-rank integrates well with the conformal setting, and experimentally validate its use in two conformal applications requiring such a correction.
Multiple testing issues in conformal prediction. We previously motivated that running multiple conformal procedures at once equates permutation testing in parallel at significance level $\alpha$ (Eq. 3). By Eq. 4 and Remark 1 this is the same as ensuring $(1-\alpha)$ coverage for an $m$-dimensional response $\mathbf{Y}_{n+1}=(Y_{n+1,1},\dots,Y_{n+1,m})$ simulatenously. Take $\mathbf{X}_{n+1}=(X_{n+1,1},\dots,X_{n+1,m})$ to be a matching multivariate input, then we directly observe how the coverage guarantee in Eq. 2 is violated since
$\mathbb{P}\left(\mathbf{Y}_{n+1}\in\hat{C}(\mathbf{X}_{n+1})\right)$ $=\mathbb{P}\left(\bigcap_{k=1}^{m}\left(Y_{n+1,k}\in\hat{C}(X_{n+1,k})\right)\right)$
$\geq 1-\mathbb{P}\left(\bigcup_{k=1}^{m}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)\right)$
$\geq 1-\sum_{k=1}^{m}\mathbb{P}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)$
$=1-\sum_{k=1}^{m}\mathbb{P}\left(\hat{P}_{n+1,k}(Y_{n+1,k};S)\leq\alpha_{k}\right)$
$\geq 1-\sum_{k=1}^{m}\alpha_{k}=1-m\alpha\quad\text{for }\alpha_{k}=\alpha\text{ fixed.}$. (7)
@@
A key insight is that due to their joint dependence on common calibration data $\mathcal{D}_{cal}$, any obtained conformal p-values are bound to exhibit some degree of positive dependency. This behaviour is referred to as positive regression dependent on a subset (PRDS) (Bates et al., 2022; Benjamini and Yekutieli, 2001; Chi et al., 2022), and highlights that any multiple testing correction relying on an independence assumption (such as Bonferroni) is bound to be overly conservative a priori.

### v4

@@ 3.1 The max-rank algorithm
Consider running $m$ hypothesis tests in parallel, and for any particular testing dimension $k\in[m]$ the statistics forming an empirical null distribution are given by $\textbf{s}_{k}\in\mathbb{R}^{n}$. In the conformal setting, $\mathbf{s}_{k}$ equates the set of nonconformity scores on $\mathcal{D}_{cal}$. We now desire finding adjusted significance levels $\hat{\alpha}_{k}$ for all dimensions such that the FWER is globally controlled at target level $\alpha$.
We algorithmically describe our proposed max-rank correction in Algo. 2 (exemplary code is given in Listing 1). First, the empirical nulls of equal size are collected in a matrix of test statistics $\mathbf{S}=[s_{i,k}]\in\mathbb{R}^{n\times m}$, where $i\in[n]$ and $k\in[m]$. The $m$ columns span the tested hypotheses, and the $n$ rows represent values in each null distribution. Next, we obtain a similar matrix containing the column-wise ranks of $\mathbf{S}$ as $\mathbf{R}=[r_{i,k}]\in\mathbb{N}^{n\times m}$, where also $i\in[n]$ and $k\in[m]$. By Eq. 4, testing at significance level $\alpha$ in each column is equivalent to computing the conformal quantile $\hat{Q}(1-\alpha;\mathbf{s}_{k})$ in the space of test statistics. This in turn equates determining the value $s_{t,k}$, where $t=\hat{Q}(1-\alpha;\mathbf{r}_{k})$ denotes a quantile in rank space ensuring the desired error control (Remark 1). The max-rank correction adjusts these rank quantiles for FWER control by forming a composite empirical null $\mathbf{r}_{\max}$ via the $l^{\infty}$-norm across tests, and using the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ to inform corrected significance levels $\hat{\alpha}_{k}$ for each hypothesis $k\in[m]$.
@@
2:Output: Adjusted quantiles with FWER control at level $\alpha$ as $(\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}})\in\mathbb{R}^{m}$.
@@
6:Compute the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ as the $\lceil(n+1)(1-\alpha)\rceil/n$-th empirical quantile of $\mathbf{r}_{\max}$.
@@
8:Select the test statistic $s_{r_{\max},k}$ at rank $r_{\max}$ column-wise in $\mathbf{S}$ as the adjusted quantile $\hat{Q}_{1-\hat{\alpha}_{k}}$ for every test dimension $k\in[m]$.
@@
Intuition. Our max-rank correction intuitively collapses multiple empirical nulls into a single ‘worst-case’ null for which the error-controlling quantile is computed. Thus, any potential dependencies across test statistics which are reflected in the rank ordering are efficiently aggregated in the ‘maximum rank’ statistic $\mathbf{r}_{\max}$, and subsequently the selected rank quantile $r_{\max}$. By virtue of applying the $l^{\infty}$-norm, $r_{\max}$ ensures that adjusted significance levels obey FWER control.
If the empirical nulls are uncorrelated and the tested hypotheses can thus be considered independent, no benefit is obtained from max-rank’s aggregation step. We prove that in such cases our approach performs at least on par with the Bonferroni correction (§ 4), and never worse. On the other hand, if the tests are perfectly correlated each empirical null exhibits identical rank ordering, and max-rank will provide the optimal quantile solution for each individual testing dimension. For any cases with positively dependent tests inbetween the two extremes, our approach ensures superior performance over Bonferroni, and scales well even as the number of hypotheses $m$ grows (see § 5). We highlight that no additional assumptions beyond those already underlying CP are required for the use of max-rank.
@@ 3.2 Relation to Westfall & Young corrections
A strong connection exists between max-rank and the max-T and min-P resampling-based MHT corrections of Westfall and Young (1993), which offer similar FWER control. The max-T method provides adjusted p-values by comparing test statistics to the distribution of maximum statistics across hypotheses (Algo. 3), while the min-P method adjusts p-values based on the minimum observed p-values after resampling (Algo. 4). In both cases, existing dependency structure between tests is leveraged through repeated data resampling.
@@
Exploring these two corrections through the lens of our conformal prediction setting, we find that max-rank can be cast as a procedure that (i) naturally extends max-T to the rank space, and (ii) describes a particular instance of min-P arising from our problem setting. As such, we argue that max-rank exhibits the advantages of both methods by permitting its direct application to obtained test statistics as in max-T, whilst maintaining scale invariance as in min-P. We next relate to each correction in more detail.
@@
Relation to max-T. Our conformal setting directly provides empirical null distributions for every hypothesis by use of the calibration set. The computed nonconformity scores on $\mathcal{D}_{cal}$ leverage available labels to produce statistics known to satisfy the tested null $H_{0}:Y_{n+1}=y$, and are thus interpretable as an analog to the required “data resampling under the null” (Westfall and Young, 1993). Similarly, CP’s exchangeability condition aligns with the exchangeability underlying the resampling step. It follows that the adjusted p-values in Algo. 3 match the conformal p-values in Eq. 3 simply by replacing the score set $S$ with a maximum score set $S_{\max}$, i.e., replacing $\hat{P}_{n+1}(y;S)$ with $\hat{P}_{n+1}(y;S_{\max})$ [^1] And correcting for $(n+1)$ samples instead of $n$.. Instead of adjusting the p-values as in Algo. 3 we may also directly target an adjustment of the conformal quantile (Eq. 4), which translates to computing $\hat{Q}(1-\alpha;S_{\max})$ instead of $\hat{Q}(1-\alpha;S)$. Comparing this to max-rank, we notice that the difference lies in its direct application to the maximum statistics, rather than translating them into rank space first (i.e., using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$). This means that max-T is sensitive to scale and requires test statistics across hypotheses to be directly comparable. Interestingly, such a “maximum of scores” approach has been intuitively employed by researchers in conformal prediction, e.g., Andéol et al. (2023); Schuster et al. (2021); Fisch et al. (2022), but never been recognized as an application case of Westfall and Young (1993).
@@
Relation to min-P. By the same resampling interpretation as above, unadjusted p-values for each hypothesis in Algo. 4 can be computed via conformal p-values (Eq. 3). Using the rank definition in Eq. 5, a given p-value $\hat{P}_{n+1}(y;S)$ can be expressed in terms of score ranking as $((n+1)-\text{rank}(s_{n+1};S))/(n+1)$ [^2] Observe that the p-value in Eq. 3 counts the fraction of values above $s_{n+1}$, whereas the rank counts values below it.. Thus, computing the minimum p-values in min-P can be interpreted as the maximum over inverted and normalized ranks. Via Eq. 4 we again translate adjusted p-values to a quantile adjustment using $\hat{Q}(\alpha;\mathbf{p}_{\min})$. Note that this quantile is now taken over p-values and only permits testing in p-value space. In contrast, max-rank keeps the simplicity of operating directly on statistics, facilitating the interpretation of testing decisions.
@@ 4 THEORETICAL RESULTS
We next briefly describe our theoretical results for max-rank, deferring details and proofs to Appendix A. The multiple testing problem is first rephrased in terms of multivariate copulas, and our two key results state that (i) max-rank provides a valid MHT correction with FWER control, and (ii) is assured to perform better than Bonferroni under positive dependency.
Recasting the multiple testing problem (§ A.1). Consider an $m$-dimensional random variable $\mathbf{X}\in\mathbb{R}^{m}$ and denote its c.d.f. as $F_{\mathbf{X}}(\mathbf{x})$. Using the interpretation in Remark 1, determining adjusted significance levels $(\hat{\alpha}_{1},\dots,\hat{\alpha}_{m})$ with joint Type-I error control at level $\alpha$ is then equivalent to finding a vector $\mathbf{x}_{*}\in\mathbb{R}^{m}$ satisfying $F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha$. By use of Sklar’s Theorem (Sklar, 1973) to express the c.d.f. in terms of a copula $C_{\mathbf{X}}$, and defining a vector of probabilities $(q_{*,1},\dots,q_{*,m})=(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))\in[0,1]^{m}$ restricted to a fixed value $q_{*}=q_{*,1}=\dots=q_{*,m}$, we can restate the problem as the search for a solution
@@
Obtaining a solution via max-rank (§ A.2). For a given solution to Eq. 6 the desired quantiles with FWER control are then computed as $\mathbf{x}_{*}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. These equate the quantiles $\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}}$ described in Algo. 2. Leveraging operations in the rank space, we show that max-rank provides such a solution $q_{\max}$ to the problem. We state this in the following proposition:
@@ 5 SIMULATION STUDY
Global FWER control. We compare the adjusted significance levels produced by both max-rank and Bonferroni across varying correlation strengths and for $m\in\{1,3,5,10,20\}$ in Fig. 1. We observe two key drawbacks for Bonferroni. Firstly, the correction becomes increasingly conservative as correlation between tests increases, with a FWER control level tending towards zero. Secondly, the rate accelerates as the number of tests increases. In contrast, max-rank suffers from neither of these behaviours and tightly maintains FWER control at target level $\alpha$, ensuring consistent results independently of correlation strength and test dimensionality (Prop. 1).
@@
FWER control and sample size (§ D.1). Since max-rank relates to Westfall and Young (1993) and is thus interpretable as a resampling-based correction, we anticipate the sample size $n$ (equivalent to the number of resamples) to affect FWER control, in particular when running many tests in parallel. Thus, we verify in Fig. 4 adjusted significance levels for max-rank and Bonferroni across varying combinations of correlation $\rho$, sample size $n$ and test counts $m$. A direct correspondence between FWER control and the magnitudes of $m$ and $n$ is apparent: as we run more tests in parallel the need for additional samples increases in order to empirically maintain FWER control at target level $\alpha$ (e.g., roughly $n\geq 100\,m$ for $\rho=0$). Crucially, larger correlations suggest a lower need for additional samples as max-rank can exploit existing rank order information, whereas Bonferroni merely grows more conservative.
@@ 6 APPLICATION TO CONFORMAL PREDICTION
We further elaborate on the usefulness of our proposed max-rank correction for conformal prediction. First, we illustrate how multiple testing issues emerge when running multiple CP procedures in parallel. We then theoretically motivate why max-rank integrates well with the conformal design and its underlying conditions, and empirically validate its effectiveness in two CP applications that require a multiple testing correction.
Multiple testing issues in conformal prediction. We previously discussed that running multiple conformal procedures at once equates permutation testing in parallel at a fixed significance level $\alpha$. By Eq. 4 and Remark 1 this is identical to ensuring a $(1-\alpha)$ coverage rate for each component of an $m$-dimensional response $\mathbf{Y}_{n+1}=(Y_{n+1,1},\dots,Y_{n+1,m})$. The same intuition is employed in our theoretical formulation for Eq. 6. Take $\mathbf{X}_{n+1}=(X_{n+1,1},\dots,X_{n+1,m})$ to be a matching multivariate input, then we can directly observe how the coverage guarantee in Eq. 2 can be violated since
$\mathbb{P}\Big{(}$ $\mathbf{Y}_{n+1}\in\hat{C}(\mathbf{X}_{n+1})\Big{)}$
$=\mathbb{P}\Big{(}\bigcap_{k=1}^{m}\big{(}Y_{n+1,k}\in\hat{C}(X_{n+1,k})\big{)}\Big{)}$
$=1-\mathbb{P}\Big{(}\bigcup_{k=1}^{m}\big{(}Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\big{)}\Big{)}$
$\geq 1-\sum_{k=1}^{m}\mathbb{P}\Big{(}Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\Big{)}$
$=1-\sum_{k=1}^{m}\mathbb{P}\Big{(}\hat{P}_{n+1,k}(Y_{n+1,k};S)\leq\alpha_{k}\Big{)}$
$\geq 1-\sum_{k=1}^{m}\alpha_{k}=1-m\alpha,\quad\text{for }\alpha_{k}=\alpha\text{ fixed.}$. (7)
@@
A key observation is that due to their joint dependence on common calibration data $\mathcal{D}_{cal}$, obtained conformal p-values in any single dimension are bound to exhibit some degree of positive dependency, a property known as positive regression dependent on a subset (PRDS) (Benjamini and Yekutieli, 2001; Bates et al., 2022; Chi et al., 2025). This highlights that any MHT correction relying on an independence assumption (such as Bonferroni) is bound to be overly conservative a priori, and our empirical results point towards a similar positive dependency across testing dimensions for CP settings.